A river flows due south with a

speed of A man steers a motorboat across the river; his

velocity relative to the water is due east. The river is 800 m

wide. (a) What is his velocity (magnitude and direction) relative to

the earth? (b) How much time is required to cross the river?

(c) How far south of his starting point will he reach the opposite

bank?


1
Expert's answer
2021-11-21T17:27:56-0500

Assume that

A river flows due south with a speed of "v_S=2\ m/s". A man steers a motorboat across the river due east. His velocity relative to the water is "v_E=4\ m/s". So,



(a) "v=\sqrt{v_E^2+v_S^2}=\sqrt{4^2+2^2}=\sqrt{20}=4.47\ (m/s)"


"\theta=\tan^{-1}(2/4)=26.6° \ \ (S \ \ of \ \ E)"


(b) "t=s/v_E=800/4=200\ (s)"


(c) "l=v_St=2\cdot200=400\ (m)"